x/5 - y/10 = 4
x/2 + y/4 = 5
Solve the system of equation using the addition method. I have tried different problems over 80+ times never getting them correct and cannot figure out what I am doing wrong. Please someone help me.
Thank you for your time. I am trying to pass my finals and this is the only problem preventing me from doing so. If I can't figure out how to do it, I cant take my finals. Help!

Answer by Edwin McCravy(20056)

Clear of fractions first:

x/5 - y/10 = 4
x/2 + y/4 = 5 

Multiply the first one through by LCD = 10
Multiply the second one through by LCD = 4

2x - y = 40
2x + y = 20
-----------
4x     = 60
     x = 15

   2x + y = 20
2(15) + y = 20
   30 + y = 20
        y = -10

(x,y) = (15,-10)

Edwin